OpenStudy (anonymous):

help!?!?! A piece of aluminum occupies a volume of 14.0 milliliters and weighs 32.6 grams. What is its density of the aluminum rounded to the nearest hundredth? Only enter numerical values, which can include a decimal point

OpenStudy (kc_kennylau):

density=mass/volume

OpenStudy (anonymous):

density=32.6/14.0 =2.328571429 right ??

OpenStudy (cggurumanjunath):

unit should be kg/m^3 or gram/cc. Common units[edit] The SI unit for density is: kilograms per cubic meter (kg/m3) Litres and metric tons are not part of the SI, but are acceptable for use with it, leading to the following units: kilograms per liter (kg/L) grams per milliliter (g/mL) metric tons per cubic meter (t/m3) Densities using the following metric units all have exactly the same numerical value, one thousandth of the value in (kg/m3). Liquid water has a density of about 1 kg/dm3, making any of these SI units numerically convenient to use as most solids and liquids have densities between 0.1 and 20 kg/dm3. kilograms per cubic decimetre (kg/dm3) grams per cubic centimetre (g/cm3) 1 gram/cm3 = 1000 kg/m3 megagrams (metric tons) per cubic metre (Mg/m3) In US customary units density can be stated in: Avoirdupois ounces per cubic inch (oz/cu in) Avoirdupois pounds per cubic inch (lb/cu in) pounds per cubic foot (lb/cu ft) pounds per cubic yard (lb/cu yd) pounds per US liquid gallon (lb/gal) pounds per US bushel (lb/bu) slugs per cubic foot Imperial units differing from the above (as the Imperial gallon and bushel differ from the US units) in practice are rarely used, though found in older documents. The density of precious metals could conceivably be based on Troy ounces and pounds, a possible cause of confusion.

OpenStudy (cggurumanjunath):

http://en.wikipedia.org/wiki/Density

OpenStudy (kc_kennylau):

@luka1998 which unit is implicitly/explicitly stated to be answered in in your book?

OpenStudy (anonymous):

the questions is not from the book its from flvs

OpenStudy (kc_kennylau):

wait, 14.0 mm can't be a unit for volume...

OpenStudy (anonymous):

but the question says that 14.0 is the volume

OpenStudy (kc_kennylau):

does it say millimeter cubed?

OpenStudy (anonymous):

no 14.0 millimeter

OpenStudy (kc_kennylau):

@hartnn

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Volume we can measure the volume from it since 1litre =0.001cubic metres

OpenStudy (anonymous):

@OOOPS is this right density=32.6/14.0 =2.328571429 ??

OpenStudy (anonymous):

@farabor help plz!

OpenStudy (kc_kennylau):

yep and the unit would be g/mm^3 but I don't know if this is the required unit

OpenStudy (anonymous):

@Euler271 can u cheek this ??

OpenStudy (anonymous):

the first answer you gave was correct. however it is very important to state the units when giving density since as demonstrated by by CGGURUMANJUNATH, there are so many different ways to represent it. g/mL is a valid density so you can keep the answer 2.33 g/mL (however, sometimes you might need to convert the units)

OpenStudy (anonymous):

@Euler271 thx

OpenStudy (anonymous):

np :)